# Gradient-Guided Annealing for Domain Generalization

### v3

@@ 1 Introduction
The disagreement of gradients across data from different domains during training, can provide an indication that the described problem occurs. This observation about gradient conflicts was initially made in the multi-task learning paradigm [92], where gradients of different tasks pointed to conflicting directions on the loss surface. To mitigate the effects of gradient conflicts, methods to balance the relative gradient magnitudes were proposed [18, 38], along with algorithms that remove disagreeing gradient components among tasks [68, 86]. Similarly, gradient conflicts were also addressed in the context of DG in [53], where gradients with different signs among domains were either muted or set to random values as they are assumed to contain domain-specific information. With that being said, it has been empirically shown that even though the above issues arise, the correct selection and tuning of hyperparameters can be enough for vanilla network training to surpass even state-of-the-art algorithms [31]. This evidence leads us to believe that there exist local minima on the vanilla loss surface that lead to more robust and generalizable models. Depending the starting parameter configuration, even a model optimized by traditional Empirical Risk Minimization can reach a solution that is locally optimal for all distinct domains. These hypotheses, along with problems poised by gradient disagreement, are demonstrated in Section 1.1 (illustrated in Fig. 1) and are further explored in Section 3.2.
Starting from these ideas, we propose an alternative strategy for updating the parameters of a neural network during the optimization process, in an attempt to “set the model up for success”. Specifically, inspired by Simulated Annealing Optimization [41], during the early stages of model training we iteratively anneal, i.e., randomly perturb the parameters of the model and search for a set of parameter values where the gradients between all training domains agree, before minimizing the total loss across all training domains. We call this simple, yet effective, strategy Gradient-Guided Annealing or GGA. When evaluated on extensive and challenging DG benchmarks, GGA is able to boost the performance of the baseline by significant margins, even yielding state-of-the-art results without the application of additional data processing or augmentation techniques. Additionally, as GGA can be considered a general strategy for handling multi-domain datasets it can also be combined with previously proposed algorithms. Experimental results of the combined methods, demonstrate the efficacy of GGA, since it is able to enhance their generalization capacity and overall accuracy, boosting their performance over the baseline benchmarks.
@@ 3.3 Gradient-Guided Annealing for DG
1:Pretrained DNN $h$ with parameters $\bm{\theta}_{0}$, training dataset with domains $\mathcal{S}=\{D_{i}\}_{1}^{\lvert\mathcal{S}\rvert}$. Loss gradients $\bm{g}$. Learning rate $\eta$, perturbation parameter $\rho$, optimization step to start the annealing process $A_{s}$, optimization step to end the annealing process $A_{e}$, number of annealing iterations per optimization step, $n_{a}$. Total number of training iterations $n$.
@@
The analysis presented in the previous section shows that we can use the agreement of gradients $\nabla\mathcal{L}_{i}$ as an indicator of domain invariance since the less the model $h(\cdot|\bm{\theta})$ depends on the domain $d$, the more similar the expected value of the gradient, $\mathbb{E}[\nabla\mathcal{L}_{d}]$ will be (see Eq. (4)). Inspired by Simulated Annealing [41], we achieve this by adding random noise to the parameters of the model in search for a point with high domain gradient agreement, as measured by the increase of the minimum gradient similarity across any pair of domains,
@@
In more detail, as it is common in DG literature we start with a pre-trained model, $f(\cdot,\bm{\theta}_{0})$ and perform a small number of warmup training steps. This ensures that the model approaches a region of the parameter space that achieves a low loss for the target problem. We then calculate the minimum pairwise gradient similarity among source domains and begin searching the neighborhood of the parameter space for points where both (a) the domain gradients agree and (b) the loss are reduced [^2] In our experiments, we relax the loss constraint and accept the new parameter set when the gradient similarity has increased and training loss has been reduced within a specific range (i.e $\mathcal{L}-\mathcal{L}^{\prime}<0.1$), where $\mathcal{L}^{\prime}$ is the loss value at the updated set of parameters.. This is implemented through iterative random perturbations $\bm{\theta}\leftarrow\bm{\theta}+\mathcal{U}(-\rho,\rho)$, where $\mathcal{U}$ is the multivariate uniform distribution. Each new point is selected only if it simultaneously achieves higher gradient agreement and lower loss (thus improving the Pareto front). After a fixed number of iterations, the point that has achieved the highest gradient agreement and the lowest loss is selected. This process is repeated after every training iteration step. After a number of optimization steps, we allow the model to be trained without parameter perturbations, following a standard SGD-based procedure, as usual. The GGA algorithm is presented in Algorithm 1.
Table 1: Comparison of GGA with the ERM baseline The top-1 out-of-domain accuracies on five domain generalization benchmarks averaged over three trials, are presented.
@@ 4.1 Experimental setup and implementation details
In all experiments, the leave-one-domain-out cross-validation protocol is followed. Specifically, in each run a single domain is left out as the target (test) domain, while the rest of the domains are used for training. The final performance of each algorithm is calculated by averaging the top-1 accuracy on the target domain, with different train-validation splits. For training, we utilize a ResNet-50 [32] pretrained on ImageNet [64] for the backbone feature extractor and ADAM for the optimizer. Regarding GGA, during training we let each algorithm run for around $50$ to $100$ training iterations depending on the dataset and then begin the parameter space search, as described in Section 3. For the neighborhood size in the search process, we set $\rho$ to 0.00050.00050.00050.0005333The selection of $\rho$ was based on previous algorithms that implement weight perturbations in ResNet-50 networks [28, 81, 45]. The sensitivity analysis presented in subsection 4.4 also reveals that $\rho=0.0005$ yields the best performance. and search for a total of $A=250$ steps before moving to the next mini-batch from each domain, in each dataset. In all experiments, we perform search iterations for $100$ to $200$ different mini-batches during early training stages. The rest of the hyperparameters, such as learning rate, weight decay and dropout rate, are tuned according to [15] and are presented in Table 3. To account for the variability introduced in the random search of GGA, we repeat the experiments with 5 different seeds for each dataset. All models were trained on a cluster containing $4\times 40$GB NVIDIA A100100100100 GPU cards, split into $8$ $20$GB virtual MIG devices and $1\times 24$GB NVIDIA RTX A5000500050005000 GPU card.
@@
Table 3: Hyperparameters for DG experiments. OH, TI and DN stand for OfficeHome, TerraIncognita and DomainNet respectively. $\rho$ is the parameter space search used during the annealing steps in GGA, while the last row is reports the training iterations during which GGA occurs.
@@ 4.2 Comparative Evaluation
The average OOD performances of the baseline vanilla ERM [74] and state-of-the-art DG methods on a total of 5 DG benchmarks, are reported in Tables 1 and 2 respectively. The results for each separate domain in each benchmark are reported in the Appendinx. In the experiments, we apply GGA on-top of the rest of the DG algorithms and show that its properties generalize to other methods as well, boosting their overall performance. It should also be noted that GGA can be adapted to training scenarios where distributions shifts are present in training data.
Initially, to validate whether the application of GGA in the early stages of model training leads to models with increased generalizabilty, we compare the results with the vanilla ERM baseline. As presented in Table 1, GGA is able to boost the performance of the baseline model by an average of 2.4% on all benchmarks, demonstrating the efficacy of the proposed algorithm. For further evaluation, we also compare GGA with state-of-the-art DG methods [49, 95, 66, 1, 91, 42, 69, 30, 34, 10, 84, 48, 71, 56, 72, 28, 98, 81, 74], before applying GGA to them as well. We note that we only include previous works that have implemented a ResNet-50 as the backbone encoder and do not use additional components or ensembles. In Table 2 we differentiate between the methods that operate on model gradients [34, 71, 28, 98, 81] and the rest of the algorithms. Even without its application to other methods, GGA is able to surpass most of the previously proposed algorithms, while also setting the state-of-the-art on PACS ($+0.7\%$) and on the challenging TerraIncognita ($+1.8$) dataset, while remaining competitive in the other two. What’s more important is that the application of GGA in conjunction with the rest of the DG methods, proves beneficial and ultimately boosts their overall performance in almost each case by an average of around $1\%$ and in some cases up to even $6.3\%$. With regards to IRM, which seems to be significantly impacted negatively by the application of GGA, its learning objective emphasizes on simultaneously minimizing the training loss of each source domain and not necessarily on the pairwise agreement of gradients among domains. It is therefore not able to converge to a good solution after the initial model’s weights have been pertrubed during training.
@@ 4.4 Sensitivity Analysis
Regarding the magnitude of weight perturbations during the application of GGA, we found that the optimal value was $\rho=5e-5$. As illustrated in Figure 4, a larger magnitude of perturbation led to decreased model performance. Intuitively, the application of larger noise to the model parameters leads to sets that are not close to solution, making it increasingly difficult for the model to converge. On the other hand, smaller perturbations seem to have little to no effect on training as the search is limited to spaces near the current parameters, which is why the model performance falls back close to that of ERM. With regards to the stage of training during which GGA will be applied, we found that the models yielded better performance when the search was initialized in earlier stages.We hypothesize that applying perturbations near the end of training displaces the model from a local optimum, requiring additional iterations to converge.
@@ 5 Conclusions
In this work, we investigated gradient conflicts in the context of domain generalization, leading to the development of GGA; a simple yet effective algorithm that identifies points in the parameter space where domain gradients align early in training before continuing optimization, ultimately enhancing model generalization. Through a comprehensive comparative analysis we have demonstrated the efficacy of GGA, which is able to achieve superior generalization capabilities in standard DG benchmarks and outperform SoTA baselines. What’s more, as GGA is both model and method agnostic it can also be utilized by other methods. Interestingly, its combination with previous algorithms proves beneficial, as their majority yields improved performance over all benchmarks. Finally, we validate the impact of our algorithm on gradient alignment through additional experiments, showing that, unlike the baseline, domain gradients align during training.
@@ F Extended Literature Review
Domain Generalization (DG) [79] is arguably one of the most difficult and fundamental problems of Machine Learning (ML) today. Unsurprisingly, a vast majority of researchers have poured effort into advancing the field, where findings have been applied to various areas, such as Natural Language Processing [35], Reinforcement Learning [48], Healthcare and Medicine [23, 4], Time-Series forecasting [25], Fault Diagnosis [93] and, of course, Computer Vision [79]. Even though not covering the entire field of DG, this section aims to present a taxonomy of the general DG methodologies developed in CV, for producing robust models that can generalize to previously unseen data, and attempts to assist potential readers navigate the past literature, while also categorizing our proposed method among its predecessors. Domain Generalization methods can be categorized into three major groups, depending on their operation during the process of model training, namely: (a) Data Manipulation, (b) Representation Learning and, (c) Learning Algorithm. Furthermore, as mentioned in the main text, DG algorithms can either leverage domain labels during training (multi-source), or completely disregard the knowledge of existing domain shifts in their training data and handle them as a single distribution (single-source).
@@
Representation Learning. This group of methods is arguably the most prominent in DG and has been the central focus of ML [9]. Following the formulation in the main text, given a labeling function $h$ that maps input observations $\bm{x}$ to their labels $y$, we can decompose it into $h=f\circ g$, where $g$ is a parametric function that learns representations of $x$ and $f$ the classifier function. The goal of representation learning can be summarized as follows:
@@
where $\ell$ the loss function to be minimized and $\ell_{\text{reg}}$ a regularizer. Methods included in this group, focus on learning an improved representation learning function $g$. The algorithms included in this group can be further categorized into three sub-groups. Feature disentanglement [88] methods intend to extract disentangled feature vectors from samples, where each dimension can be linked to a subset of data generating factors. The main idea is to produce a model that extracts a representation that can be further decomposed into domain-specific, domain-invariant, and class-specific features. To that end, the authors of [61] present CSD, which jointly learns a domain-invariant and domain-specific component in the final representation and enables the extraction of disentangled representations, whereas the authors of [16] propose learning domain specific masks during training to improve model robustness. Generative models have also been proposed in the disentangled representation learning literature for DG, with variational autoencoders (VAEs) and GANs [17] being utilized for learning distinct latent subspaces for class- and domain-specific features [36]. Another promising category of methods aiming to produce disentangled representations is that of Causally-Inspired algorithms. In causal representation learning, a domain shift can be thought of as an intervention and models aim to uncover the true causal data generating factors. Naturally, the prediction of a model should not be affected by interventions on spuriously correlated but irrelevant features, such as the background, color or style of the image. Under this causal consideration, the authors of [52] propose a structural causal model in order to model within-class variations and leverage the fact that inputs across domains should have the same representation, given that they derive from the same object. Similar to disentangled representations, there have been proposed methods in the literature that focus on completely disregarding domain-relevant from the final feature vectors, deriving solely domain-invariant representations. Based on the initial findings of [8], numerous works have presented algorithms that aim to minimize the representation differences across multiple source domains within a specific feature space, ensuring they become domain invariant, ultimately enabling the trained model to effectively generalize to previously unseen domains. In one of the most notable previous works in this category, Arjovsky et al. [1] enforce the optimal classifier on top of the representation space to be the same across domains and simultaneously minimize the loss across distributions. The above idea of Invariant Risk Minimization (IRM) has been extended to several other works. For example, the authors of [43] propose minimizing the variance of source-domain risks, by minimizing their extrapolated risk, while the authors of [91] propose adapting to domain shift and producing invariant representations. Finally, an alternative route towards learning generalized representations is via regularization strategies. The most representative group of methods in this category is Gradient-Based operations, which utilize gradient information during model training. In [34], the authors propose learning robust representations by discarding the most dominant gradients in each training iteration under the assumption that they are correlated with domain-specific features present in the source data. Another popular strategy is to seek for flat minima [28, 15] in the loss landscape of neural networks during training, assuming that models that converge to flat minima exhibit increased generalization capabilities [98, 81]. What’s more, Shi et al. [71] hypothesized that gradients among domains should match and proposed an approximation of a loss inducing the maximization of the gradient inner product during training. Our method (GGA) can be categorized in this group of gradient operations, as it considers the similarity of domain gradients in the early iterations of model training and seeks for sets in the parameter space with increased gradient alignment, before continuing the optimization procedure.

### v4

@@ 1 Introduction
The disagreement of gradients across data from different domains during training, can provide an indication that the described problem occurs. This observation about gradient conflicts was initially made in the multi-task learning paradigm [92], where gradients of different tasks pointed to conflicting directions on the loss surface. To mitigate the effects of gradient conflicts, methods to balance the relative gradient magnitudes were proposed [18, 38], along with algorithms that remove disagreeing gradient components among tasks [68, 86]. Similarly, gradient conflicts were also addressed in the context of DG in [53], where gradients with different signs among domains were either muted or set to random values; as they are assumed to contain domain-specific information. With that being said, it has been empirically shown that even though the above issues arise, the correct selection and tuning of hyperparameters can be enough for vanilla network training to surpass even state-of-the-art algorithms [31]. This evidence leads us to believe that there exist local minima on the vanilla loss surface that lead to more robust and generalizable models. Depending on the starting parameter configuration, even a model optimized by traditional Empirical Risk Minimization can reach a solution that is locally optimal for all distinct domains. These hypotheses, along with problems poised by gradient disagreement, are demonstrated in Section 1.1 (illustrated in Fig. 1) and are further explored in Section 3.2.
Starting from these ideas, we propose an alternative strategy for updating the parameters of a neural network during the optimization process, in an attempt to “set the model up for success”. Specifically, inspired by Simulated Annealing Optimization [41], during the early stages of model training we iteratively anneal, i.e. randomly perturb the parameters of the model, and search for a set of parameter values where the gradients between all training domains agree, before minimizing the total loss across all training domains. We call this simple, yet effective, strategy Gradient-Guided Annealing or GGA. When evaluated on extensive and challenging DG benchmarks, GGA is able to boost the performance of the baseline by significant margins, even yielding state-of-the-art results without the application of additional data processing or augmentation techniques. Additionally, since GGA can be considered a general strategy for handling multi-domain datasets it can also be combined with previously proposed algorithms. Experimental results of the combined methods, demonstrate the efficacy of GGA, as it is able to enhance their generalization capacity and overall accuracy, boosting their performance over the baseline benchmarks.
@@ 3.3 Gradient-Guided Annealing for DG
1:Pretrained DNN $h$ with parameters $\bm{\theta}_{0}$, training dataset with domains $\mathcal{S}=\{D_{i}\}_{1}^{\lvert\mathcal{S}\rvert}$. Loss gradients $\bm{g}$. Learning rate $\eta$, perturbation parameter $\rho$, optimization step to start the annealing process $A_{s}$, optimization step to end the annealing process $A_{e}$, number of annealing iterations per optimization step $n_{a}$. Total number of training iterations $n$.
@@
The analysis presented in the previous section shows that we can use the agreement of gradients $\nabla\mathcal{L}_{i}$ as an indicator of domain invariance, since the less the model $h(\cdot|\bm{\theta})$ depends on the domain $d$, the more similar the expected value of the gradient, $\mathbb{E}[\nabla\mathcal{L}_{d}]$, will be (see Eq. (4)). Inspired by Simulated Annealing [41], we achieve this by adding random noise to the parameters of the model in search for a point with high domain gradient agreement, as measured by the increase of the minimum gradient similarity across any pair of domains,
@@
In more detail, as it is common in DG literature, we start with a pre-trained model $f(\cdot,\bm{\theta}_{0})$ and perform a small number of warmup training steps. This ensures that the model approaches a region of the parameter space that achieves a low loss for the target problem. We then calculate the minimum pairwise gradient similarity among source domains and begin searching the neighborhood of the parameter space for points where both: (a) the domain gradients agree and (b) the loss are reduced [^2] In our experiments, we relax the loss constraint and accept the new parameter set when the gradient similarity has increased and training loss has been reduced within a specific range (i.e $\mathcal{L}-\mathcal{L}^{\prime}<0.1$), where $\mathcal{L}^{\prime}$ is the loss value at the updated set of parameters.. This is implemented through iterative random perturbations $\bm{\theta}^{\prime}\leftarrow\bm{\theta}+\mathcal{U}(-\rho,\rho)$, where $\mathcal{U}$ is the multivariate uniform distribution. Each new point is selected only if it simultaneously achieves higher gradient agreement and lower loss (thus improving the Pareto front). After a fixed number of iterations, the point that has achieved the highest gradient agreement and the lowest loss is selected. This process is repeated after every training iteration step. After a number of optimization steps, we allow the model to be trained without parameter perturbations, following a standard SGD-based procedure, as usual. The GGA algorithm is presented in Algorithm 1.
Table 1: Comparison of GGA with the ERM baseline. The top out-of-domain accuracies on five domain generalization benchmarks averaged over three trials, are presented.
@@ 4.1 Experimental setup and implementation details
In all experiments, the leave-one-domain-out cross-validation protocol is followed. Specifically, in each run a single domain is left out as the target (test) domain, while the rest of the domains are used for training. The final performance of each algorithm is calculated by averaging the top-1 accuracy on the target domain, with different train-validation splits. For training, we utilize a ResNet-50 [32] pretrained on ImageNet [64] for the backbone feature extractor and ADAM for the optimizer. Regarding GGA, during training we let each algorithm run for several training iterations depending on the dataset and then begin the parameter space search, as described in Section 3. For the neighborhood size in the search process, we set $\rho$ to 0.00050.00050.00050.0005333The selection of $\rho$ was based on previous algorithms that implement weight perturbations in ResNet-50 networks [28, 81, 45]. The sensitivity analysis presented in subsection 4.4 also reveals that $\rho=0.0005$ yields the best performance. and search for a total of $A=250$ steps before moving to the next mini-batch of 48 samples from each domain, in each dataset. In all experiments, we perform search iterations for $100$ different mini-batches during early training stages. The rest of the hyperparameters, such as learning rate, weight decay and dropout rate, are tuned according to [15] and are presented in Table 3. To account for the variability introduced in the random search of GGA, we repeat the experiments with 5 different seeds for each dataset. All models were trained on a cluster containing $4\times 40$GB NVIDIA A100100100100 GPU cards, split into $8$ $20$GB virtual MIG devices and $1\times 24$GB NVIDIA RTX A5000500050005000 GPU card.
@@
Table 3: Hyperparameters for DG experiments. OH, TI and DN stand for OfficeHome, TerraIncognita and DomainNet respectively. $\rho$ is the parameter space search used during the annealing steps in GGA, while the last row indicates the training iterations during which GGA occurs.
@@ 4.2 Comparative Evaluation
The average OOD performances of the baseline vanilla ERM [74] and state-of-the-art DG methods on a total of 5 DG benchmarks, are reported in Tables 1 and 2 respectively. The results for each separate domain in each benchmark are reported in the Appendix. In the experiments, we apply GGA on-top of the rest of the DG algorithms and show that its properties generalize to other methods as well, boosting their overall performance. It should also be noted that GGA can be adapted to training scenarios where distributions shifts are present in training data.
Initially, to validate whether the application of GGA in the early stages of model training leads to models with increased generalizabilty, we compare the results with the vanilla ERM baseline. As presented in Table 1, GGA is able to boost the performance of the baseline model by an average of 2.4% on all benchmarks, demonstrating the efficacy of the proposed algorithm. For further evaluation, we also compare GGA with state-of-the-art DG methods [49, 95, 66, 1, 91, 42, 69, 30, 34, 10, 84, 48, 71, 56, 72, 28, 98, 81, 74], before applying GGA to them as well. We note that we only include previous works that have implemented a ResNet-50 as the backbone encoder and do not use additional components or ensembles. In Table 2 we differentiate between the methods that operate on model gradients [34, 71, 28, 98, 81] and the rest of the algorithms. Even without its application to other methods, GGA is able to surpass most of the previously proposed algorithms, while also setting the state-of-the-art on PACS ($+0.7\%$) and on the challenging TerraIncognita ($+1.8$) dataset, while remaining competitive in the other two. What’s more important is that the application of GGA in conjunction with the rest of the DG methods, proves beneficial and ultimately boosts their overall performance in almost each case by an average of around $1\%$ and in some cases up to even $6.3\%$. With regards to IRM, which seems to be significantly impacted negatively by the application of GGA, its learning objective emphasizes on simultaneously minimizing the training loss of each source domain and not necessarily on the pairwise agreement of gradients among domains. It is therefore not able to converge to a good solution after the initial model’s weights have been perturbed during training.
@@ 4.4 Sensitivity Analysis
Regarding the magnitude of weight perturbations during the application of GGA, we found that the optimal value was $\rho=5e-5$. As illustrated in Figure 4, a larger magnitude of perturbation led to decreased model performance. Intuitively, the application of larger noise to the model parameters leads to sets that are not close to the solution, making it increasingly difficult for the model to converge. On the other hand, smaller perturbations seem to have little to no effect on training, as the search is limited to spaces near the current parameters, which is why the model performance falls back close to that of ERM. With regards to the stage of training during which GGA will be applied, we found that the models yielded better performance when the search was initialized in earlier stages.We hypothesize that applying perturbations near the end of training displaces the model from a local optimum, requiring additional iterations to converge.
@@ 5 Conclusions
In this work, we investigated gradient conflicts in the context of domain generalization, leading to the development of GGA; a simple yet effective algorithm that identifies points in the parameter space where domain gradients align early in training before continuing optimization, ultimately enhancing model generalization. Through a comprehensive comparative analysis we have demonstrated the efficacy of GGA, which is able to achieve superior generalization capabilities in standard DG benchmarks and outperform SoTA baselines. What’s more, as GGA is both model and method agnostic it can also be utilized by other algorithms. Interestingly, its combination with previous methods proves beneficial, as their majority yields improved performance over all benchmarks. Finally, we validate the impact of our algorithm on gradient alignment through additional experiments, showing that, unlike the baseline, domain gradients align during training.
@@ F Extended Literature Review
Domain Generalization (DG) [79] is arguably one of the most difficult and fundamental problems of Machine Learning (ML) today. Unsurprisingly, a vast number of researchers have poured effort into advancing the field, where findings have been applied to various areas, such as Natural Language Processing [35], Reinforcement Learning [48], Healthcare and Medicine [23, 4], Time-Series forecasting [25], Fault Diagnosis [93] and, of course, Computer Vision [79]. Even though not covering the entire field of DG, this section aims to present a taxonomy of the general DG methodologies developed in CV, for producing robust models that can generalize to previously unseen data, and attempts to assist potential readers navigate the past literature, while also categorizing our proposed method among its predecessors. Domain Generalization methods can be categorized into three major groups, depending on their operation during the process of model training, namely: (a) Data Manipulation, (b) Representation Learning and, (c) Learning Algorithm. Furthermore, as mentioned in the main text, DG algorithms can either leverage domain labels during training (multi-source), or completely disregard the knowledge of existing domain shifts in their training data and handle them as a single distribution (single-source).
@@
Representation Learning. This group of methods is arguably the most prominent in DG and has been the central focus of ML [9]. Following the formulation in the main text, given a labeling function $h$ that maps input observations $\bm{x}$ to their labels $y$, we can decompose it into $h=f\circ g$, where $g$ is a parametric function that learns representations of $x$ and $f$ is the classifier function. The goal of representation learning can be summarized as follows:
@@
where $\ell$ the loss function to be minimized and $\ell_{\text{reg}}$ a regularizer. Methods included in this group, focus on learning a robust and generalizable representation learning function $g$. The algorithms included in this group can be further categorized into three sub-groups. Feature disentanglement [88] methods intend to extract disentangled feature vectors from samples, where each dimension can be linked to a subset of data generating factors. The main idea is to produce a model that extracts a representation that can be further decomposed into domain-specific, domain-invariant, and class-specific features. To that end, the authors of [61] present CSD, which jointly learns a domain-invariant and domain-specific component in the final embedding and enables the extraction of disentangled representations, whereas the authors of [16] propose learning domain specific masks during training to improve model robustness. Generative models have also been proposed in the disentangled representation learning literature for DG, with variational autoencoders (VAEs) and GANs [17] being utilized for learning distinct latent subspaces for class- and domain-specific features [36]. Another promising category of methods aiming to produce disentangled representations is that of Causality-Inspired algorithms. In causal representation learning, a domain shift can be thought of as an intervention, subsequentially leading the development of models that aim to uncover the true causal data generating factors. Naturally, the prediction of a model should not be affected by interventions on spuriously correlated but irrelevant features, such as the background, color or style of the image. Under this causal consideration, the authors of [52] propose a structural causal model in order to model within-class variations and leverage the fact that inputs across domains should have the same representation, given that they derive from the same object. Similar to disentangled representations, there have been proposed methods in the literature that focus on completely disregarding domain-relevant from the final feature vectors, deriving solely domain-invariant representations. Based on the initial findings of [8], numerous works have presented algorithms that aim to minimize the representation differences across multiple source domains within a specific feature space, ensuring they become domain invariant, ultimately enabling the trained model to effectively generalize to previously unseen domains. In one of the most notable previous works in this category, Arjovsky et al. [1] enforce the optimal classifier on top of the representation space to be the same across domains and simultaneously minimize the loss across distributions. The above idea of Invariant Risk Minimization (IRM) has been extended to several other works. For example, the authors of [43] propose minimizing the variance of source-domain risks, by minimizing their extrapolated risk, while the authors of [91] propose adapting to domain shift and producing invariant representations. Finally, an alternative route towards learning generalized representations is via regularization strategies. The most representative group of methods in this category is Gradient-Based operations, which utilize gradient information during model training. In [34], the authors propose learning robust representations by discarding the most dominant gradients in each training iteration under the assumption that they are correlated with domain-specific features present in the source data. Another popular strategy is to seek for flat minima [28, 15] in the loss landscape of neural networks during training, assuming that models that converge to flat minima exhibit increased generalization capabilities [98, 81]. What’s more, Shi et al. [71] hypothesized that gradients among domains should match and proposed an approximation of a loss inducing the maximization of the gradient inner product during training. Our method (GGA) can be categorized in this group of gradient operations, as it considers the similarity of domain gradients in the early iterations of model training and seeks for sets in the parameter space with increased gradient alignment, before continuing the optimization procedure.